OpenStudy (anonymous):

What is the solution to the following equation? 5(2x − 6) + 20 = 10 9 5 3 2

pooja195 (pooja195):

5(2x − 6) Distribute

OpenStudy (anonymous):

well i guess i should start off by saying i got 2 as my asnwer

pooja195 (pooja195):

Thats correct

OpenStudy (anonymous):

oh ok thanks :D

pooja195 (pooja195):

5(2x−6)+20=10 Step 1 5(2x−6)+20=10 Simplify 5(2x−6)+20=10 (5)(2x)+(5)(−6)+20=10(Distribute) 10x+−30+20=10 (10x)+(−30+20)=10(Combine Like Terms) 10x+−10=10 10x−10=10 Step 2: Add 10 to both sides. 10x−10+10=10+10 10x=20 Step 3: Divide both sides by 10. 10x 10 = 20/10 x=2

OpenStudy (campbell_st):

wow I would have subtracted 20 as the 1st step \[5(2x - 6) + 20 -20 = 10 -20 ~~~so ~~~~ 5(2x -6) = -10\] then divide both sides by 5 \[\frac{5(2x -6)}{5} = \frac{-10}{5} ~~~~so~~~~ 2x - 6 = -2\] next add 6 to both sides of the equation \[2x - 6 + 6 = -2 + 6~~~~or~~~~2x = 4\] now solve for x it saves a few steps

OpenStudy (abb0t):

x = 2